Phil and Inez start from the same place at the same time and walk in opposite directions on Tremont Street. Inez walks twice as fast as Phil. After 30 minutes they are 24 blocks apart. How far does Phil walk?
Answer by DrBeeee(684)

Let p = rate at which Phil walks
Let i = rate at which Inez walks
Givens:
(1) i = 2p
(2) distance walked by Phil PLUS the distance walked by Inez is 24 blocks
(3) the that they walked is 30 minutes or one half an hour
Use the distance formula
(4) d = rt in (2) to get
(5) pt + 2pt = 24
Now let t = 1/2 hour to get
(6) p/2 + 2p/2 = 24 or
(7) 3p = 48 or
(8) p = 16 blocks/hour
Then from (1) we have
(9) i = 2p or
(10) i = 32 blocks/hour
Check this with (5).
Is (16/2 + 32/2 = 24)?
Is (8 + 16 = 24)?
Is (24 = 24)? Yes
Using (1) we can find out how far Phil walked
(11) Phil walked = pt or
(12) Phil walked = 16*1/2 or
(13) Phil walked = 8
Answer: Phil walked 8 blocks (Inez walked 16 blocks).